# LLM one-shot style transfer for Authorship Attribution and Verification

### v1

@@ Abstract
Computational stylometry analyzes writing style through quantitative patterns in text, supporting applications from forensic tasks such as identity linking and plagiarism detection to literary attribution in the humanities. Supervised and contrastive approaches rely on data with spurious correlations and often confuse style with topic. Despite their natural use in AI-generated text detection, the CLM pre-training of modern LLMs has been scarcely leveraged for general authorship problems. We propose a novel unsupervised approach based on this extensive pre-training and the in-context learning capabilities of LLMs, employing the log-probabilities of an LLM to measure style transferability from one text to another. Our method significantly outperforms LLM prompting approaches of comparable scale and achieves higher accuracy than contrastively trained baselines when controlling for topical correlations. Moreover, performance scales fairly consistently with the size of the base model and, in the case of authorship verification, with an additional mechanism that increases test-time computation; enabling flexible trade-offs between computational cost and accuracy.
Keywords Zero-shot $\cdot$ Authorship Attribution $\cdot$ Authorship Verification
@@ 1 Introduction
Computational stylometry is the quantitative and computational analysis of writing style, utilizing patterns in vocabulary, syntax, punctuation, and character sequences to identify, verify, or cluster authors based on textual data [1]. Authorship analysis has been widely applied across diverse domains, with applications spanning both forensic and humanities contexts. In forensics, stylistic and linguistic patterns have been leveraged to analyze terrorist discourse [2], link user identities across online platforms [3], detect plagiarism [4], and identify coordinated disinformation campaigns [5]. Beyond these security-driven settings, authorship methods also play a key role in the humanities, supporting tasks such as the attribution of pseudepigraphic texts [6] and the analysis of anonymous literary works [7].
Authorship analysis is a long-standing challenge in natural language processing, with early approaches that take advantage of hand-crafted stylometric features [8, 9]. The advent of pre-trained transformers [10, 11] led to approaches using pre-trained encoder-only transformers [12, 13], allowing more sophisticated encoding of semantic information, yet these models often struggle in the absence of topical correlations [13].
A different supervised approach in the literature involves fine-tuning individual language models for candidate authors [14, 15], measuring the probability assigned to texts by the language model. While this approach tries to leverage the pre-training of LLMs, it is expensive to apply, and not useful in low resource contexts like authorship verification tasks.
In this work, we focus on non-supervised approaches, given the tendency of supervised training to produce models biased towards semantic similarity. In this category, we include semi-supervised contrastive learning approaches [16, 17, 18], which still remain limited by the reliance on labeled data. The emergence of large language models (LLMs) has introduced prompt-based unsupervised stylistic analysis [19, 20], offering an alternative without supervision but facing poor performance (at least at moderate model sizes; see fig.˜5) and challenges in attribution due to context length constraints.
@@ 2.2 Available data
Authorship datasets are foundational for advancing research in computational authorship analysis. Diverse corpora exist across multiple domains, including news articles [21], opinion pieces [22], product reviews [23], blogs [24], literary fiction [25], and social media platforms such as Reddit [26] and Twitter [27]. For robust evaluation, this work focuses on datasets that have been cleaned and standardized for AA and AV tasks.
A significant contribution to dataset quality and diversity comes from the CLEF Initiative’s annual PAN competitions, which set benchmarks for digital text forensics and stylometry. Early AA datasets were introduced in 2011 [28] and 2012 [29], drawing from corporate emails and fiction, respectively. The 2018 PAN AA task [30, 31] established a challenging cross-fandom scenario within fanfiction, where all unknown documents originate from a single fandom, while candidate authors’ documents span non-overlapping fandoms. This design amplifies complexity by introducing domain shift and reducing stylistic overlap, as authors often emulate the source material. The 2019 task [32, 33] further advanced these challenges through an open-set evaluation.
PAN competitions have similarly contributed standardized AV datasets. The 2020 and 2021 editions [34, 35] centered on fanfiction, with the latter introducing an out-of-distribution test set featuring unseen authors and novel topics. In 2022 [36], the AV task expanded to encompass diverse discourse types, including essays, emails, text messages, and business memos—highlighting the impact of medium on writing style and increasing task difficulty.
Additionally, we leverage AV datasets derived from the PAN style-change detection tasks. These require models to determine whether consecutive paragraphs are authored by the same individual. We consider data from the 2022 [37], 2023 [38], and 2024 [39] competitions. The 2022 dataset features StackExchange posts, while the 2023 and 2024 datasets focus on Reddit, specifically curating subsets where input texts revolve around the same topic. This setting limits reliance on semantic cues and accentuates the need for nuanced stylistic detection.
@@ 2.3 Authorship methods
Traditional authorship analysis methods relied on handcrafted lexical (e.g., character and word n-grams, word frequency, word-length distribution), syntactic (e.g., POS distributions, syntactic n-grams), and semantic features [8, 9]. These features were typically used in supervised settings with classifiers such as SVMs [8], or with early neural architectures including RNNs [40] and CNNs [41].
With the introduction of transformers [10], pre-trained encoder-only models such as BERT [11] have been widely adopted for authorship-related tasks [12, 13]. These supervised models primarily capture semantic features, which limits their effectiveness in settings where texts share a common topic [13].
An alternative line of work fine-tunes an individual language model for each candidate author using a corpus of known texts, and attributes authorship by evaluating perplexity under these author-specific models [14, 15]. However, the need to fine-tune a language model for each author makes these methods expensive and not applicable in low-resource environments like authorship verification tasks.
More recently, there has been growing interest in learning author embeddings, where each text is mapped into a vector representation capturing stylistic properties. These methods often rely on contrastive learning [16, 17, 18] and assess authorship via similarity in the embedding space. However, such approaches remain dependent on labeled data that is often confounded with topical similarity.
The latest trend explores the use of large language models (LLMs) for human-like stylistic analysis based on prompts [19, 20]. Although promising because of their unsupervised nature, these methods require very large models in attribution settings because of context limitations, and empirical evaluation remains scarce. Our results indicate that, in practice, such approaches yield very limited performance in authorship verification.
@@ 2.4 Causal Language Modeling and few-shot learning
Causal Language Modeling (CLM) refers to conditioning on prior tokens to predict the next token in a unidirectional, autoregressive manner, ie, modeling $P(x_{t}\mid x_{<t})$ [42]. It is the pre-training objective used by GPT-style decoder-only models [43]. A landmark discovery in large language models (LLMs) is their few‑shot in‑context learning ability [44]. GPT 3 (175B parameters) was shown to perform new tasks with just a few textual examples provided in the prompt, without any gradient updates or fine-tuning, demonstrating competitive performance with fully fine-tuned systems [45]. This emergent capability highlights that massive autoregressive pre-training endows models with flexible, task‑agnostic generalization using only prompt examples.
@@ 4.1 Authorship attribution
We evaluated our models in established authorship attribution benchmarks in the fiction domain, including PAN12 [29], PAN18 [30], and PAN19 [32], and in the corporate email domain with PAN11 [28]. Our main baseline is STAR [18], the strongest contrastive learning method reported in prior work. Additionally, we do not compare against LLM prompting approaches such as LIP [20], as these methods require large context windows in AA, resulting in markedly lower performance at moderate model sizes.
@@ 4.2 Authorship verification
For authorship verification (AV), we evaluate on the 2020 and 2021 editions of the PAN AV shared task [34, 35], which focus on fanfiction. In addition, we construct AV datasets from the PAN style-change detection tasks of 2022–2024 [37, 38, 39]. The 2022 edition contains StackExchange posts, while the 2023 and 2024 editions comprise Reddit posts and explicitly control for topic, thereby preventing detectors from exploiting semantic information.
As in authorship attribution (AA), we compare against the STAR baseline [18]. Given that AV requires less context in LLM prompts than AA, we additionally include LIP [20] and PromptAV [19] with a 24B Mistral LLM as zero-shot baselines. For STAR and OSST-based methods, we select a decision boundary with 2000 examples from the training set.
@@
The results on authorship verification datasets are summarized in figs.˜4 and 5. Several key patterns emerge: (i) Performance on challenging tasks:nOSST consistently surpasses STAR on the more difficult settings (i.e., the third tasks in the style change benchmarks), even in its zero-shot (ZS) configuration with an 8B model. In contrast, STAR maintains an advantage in datasets with strong topic biases. (ii) Comparison to LLM prompting:At comparable model scales (up to 24B parameters), nOSST methods substantially outperform LLM-based prompting approaches. (iii) Impact of the RS variant:The RS variant of nOSST generally improves performance across multiple datasets, though a few exceptions remain. (iv) Scaling behavior:Performance exhibits a clear scaling trend—larger nOSST base models tend to achieve higher accuracy, with only minor deviations from monotonic improvement.
@@ 5 Conclusions
In this work, we propose a novel unsupervised approach to the AA and AV problems. Our OSST-based methods leverage the extensive CLM pre-training of modern LLMs, allowing greater scaling without labeled authorship data, often containing topical correlations. Our experiments suggest that OSST-based methods are very superior to LLM prompting alternatives at similar scale and perform better than contrastively trained baselines where topical correlations are controlled. In addition, they seem to confirm that performance improves with base model scale and, in the case of nOSST-RS for verification, with the number of anchor examples up to a point. This allows the user to trade-off computational resources for accuracy.
Our work presents two main limitations. First, the scale of base models used in OSST-based methods is limited due to resource constraints. More study is needed to observe scaling laws in performance. Second, in the RS variation for authorship verification, we pick in-distribution anchors as the baseline. Although this variation did increase performance, sampling anchors in-distribution in the wild is not simple. Models trained with contrastive learning could find good use for this task.
@@ Appendix A Model and prompt to obtain a neutrally styled version of a text
We use a google/gemma-3-12b-it [46] model to obtain neutral versions of a given text. This decision was made for its multilingual capabilities and the inclusion of multilingual texts in some of the datasets used.
@@ Appendix D Base models used in OSST
  • google/gemma-3-1b-pt [46]
  • google/gemma-3-4b-pt [46]
  • meta-llama/Llama-3.2-1B [47]
  • meta-llama/Llama-3.2-8B [47]

### v2

@@ Abstract
Computational stylometry studies writing style through quantitative textual patterns, enabling applications such as authorship attribution, identity linking, and plagiarism detection. Existing supervised and contrastive approaches often rely on datasets with spurious correlations, conflating style with topic. Despite the relevance of language modeling to these tasks, the pre-training of modern large language models (LLMs) has been underutilized in general authorship analysis. We introduce an unsupervised LLM-based framework that leverages CLM pre-training and in-context learning to measure style transferability between texts using log-probabilities. This approach avoids explicit supervision and topic bias while exploiting the rich stylistic knowledge encoded in LLMs. Our method substantially outperforms unsupervised prompting-based baselines of similar scale and exceeds contrastively trained models when controlling for topical overlap. Performance scales consistently with model size and, in the case of authorship verification, with an additional mechanism that increases test-time computation; enabling flexible trade-offs between computational cost and accuracy.
Keywords Unsupervised $\cdot$ Authorship Attribution $\cdot$ Authorship Verification
@@ 1 Introduction
Computational stylometry is the quantitative and computational analysis of writing style, utilizing patterns in vocabulary, syntax, punctuation, and character sequences to identify, verify, or cluster authors based on textual data [daelemans2013ExplanationComputational]. Authorship analysis has been widely applied across diverse domains, with applications spanning both forensic and humanities contexts. In forensics, stylistic and linguistic patterns have been leveraged to analyze terrorist discourse [abbasi2005ApplyingAuthorship], link user identities across online platforms [shu2017UserIdentity], detect plagiarism [stamatatos2011Plagiarismauthorship], and identify coordinated disinformation campaigns [stiff2022Detectingcomputergenerated]. Beyond these security-driven settings, authorship methods also play a key role in the humanities, supporting tasks such as the attribution of pseudepigraphic texts [koppel2013Automaticallyidentifying] and the analysis of anonymous literary works [stover2016Computationalauthorship].
Authorship analysis is a long-standing challenge in natural language processing, with early approaches that take advantage of hand-crafted stylometric features [stamatatos2016AuthorshipVerification, lagutina2019SurveyStylometric]. The advent of pre-trained transformers [vaswani2017AttentionAll, devlin2019BERTPretraining] led to approaches using pre-trained encoder-only transformers [fabien-etal-2020-bertaa, zangerle2023Overviewmultiauthor], allowing more sophisticated encoding of semantic information, yet these models often struggle in the absence of topical correlations [zangerle2023Overviewmultiauthor].
A different supervised approach in the literature involves fine-tuning individual language models for candidate authors [huang2025Attributingauthorship, barlas2020CrossDomainAuthorship], measuring the probability assigned to texts by the language model. While this approach tries to leverage the pre-training of LLMs, it is expensive to apply, and not useful in low resource contexts like authorship verification tasks.
In this work, we focus on non-supervised approaches, given the tendency of supervised training to produce models biased towards semantic similarity. In this category, we include semi-supervised contrastive learning approaches [ai2022WhodunitLearning, huertas-tato2024PARTPretrained, huertas-tato2024Understandingwriting], which still remain limited by the reliance on labeled data. The emergence of large language models (LLMs) has introduced prompt-based unsupervised stylistic analysis [hung2023WhoWrote, huang2024CanLarge], offering an alternative without supervision but facing poor performance (at least at moderate model sizes; see fig.˜5) and challenges in attribution due to context length constraints.
@@ 2.2 Available data
Authorship datasets are foundational for advancing research in computational authorship analysis. Diverse corpora exist across multiple domains, including news articles [lewis2004RCV1new], opinion pieces [stamatatos2013RobustnessAuthorship], product reviews [seroussi2014AuthorshipAttribution], blogs [schler2006Effectsage], literary fiction [gerlach2020StandardizedProject], and social media platforms such as Reddit [volske2017TLDRMininga] and Twitter [cheng2010Youare]. For robust evaluation, this work focuses on datasets that have been cleaned and standardized for AA and AV tasks.
A significant contribution to dataset quality and diversity comes from the CLEF Initiative’s annual PAN competitions, which set benchmarks for digital text forensics and stylometry. Early AA datasets were introduced in 2011 [argamon_2011_3713246] and 2012 [juola_2012_3713273], drawing from corporate emails and fiction, respectively. The 2018 PAN AA task [kestemont_2018_3737849, kestemont2018overview] established a challenging cross-fandom scenario within fanfiction, where all unknown documents originate from a single fandom, while candidate authors’ documents span non-overlapping fandoms. This design amplifies complexity by introducing domain shift and reducing stylistic overlap, as authors often emulate the source material. The 2019 task [kestemont_2019_3530313, kestemont2019overview] further advanced these challenges through an open-set evaluation.
PAN competitions have similarly contributed standardized AV datasets. The 2020 and 2021 editions [bischoff2020ImportanceSuppressing, bevendorff2020PAN20Authorship] centered on fanfiction, with the latter introducing an out-of-distribution test set featuring unseen authors and novel topics. In 2022 [stamatatos2022PAN22Authorship], the AV task expanded to encompass diverse discourse types, including essays, emails, text messages, and business memos—highlighting the impact of medium on writing style and increasing task difficulty.
Additionally, we leverage AV datasets derived from the PAN style-change detection tasks. These require models to determine whether consecutive paragraphs are authored by the same individual. We consider data from the 2022 [zangerle2022PAN22Authorship], 2023 [zangerle2023PAN23MultiAuthor], and 2024 [zangerle2024PAN24MultiAuthor] competitions. The 2022 dataset features StackExchange posts, while the 2023 and 2024 datasets focus on Reddit, specifically curating subsets where input texts revolve around the same topic. This setting limits reliance on semantic cues and accentuates the need for nuanced stylistic detection.
@@ 2.3 Authorship methods
Traditional authorship analysis methods relied on handcrafted lexical (e.g., character and word n-grams, word frequency, word-length distribution), syntactic (e.g., POS distributions, syntactic n-grams), and semantic features [stamatatos2016AuthorshipVerification, lagutina2019SurveyStylometric]. These features were typically used in supervised settings with classifiers such as SVMs [stamatatos2016AuthorshipVerification], or with early neural architectures including RNNs [bagnall2016AuthorIdentification] and CNNs [shrestha-etal-2017-convolutional].
With the introduction of transformers [vaswani2017AttentionAll], pre-trained encoder-only models such as BERT [devlin2019BERTPretraining] have been widely adopted for authorship-related tasks [fabien-etal-2020-bertaa, zangerle2023Overviewmultiauthor]. These supervised models primarily capture semantic features, which limits their effectiveness in settings where texts share a common topic [zangerle2023Overviewmultiauthor].
An alternative line of work fine-tunes an individual language model for each candidate author using a corpus of known texts, and attributes authorship by evaluating perplexity under these author-specific models [huang2025Attributingauthorship, barlas2020CrossDomainAuthorship]. However, the need to fine-tune a language model for each author makes these methods expensive and not applicable in low-resource environments like authorship verification tasks.
More recently, there has been growing interest in learning author embeddings, where each text is mapped into a vector representation capturing stylistic properties. These methods often rely on contrastive learning [ai2022WhodunitLearning, huertas-tato2024PARTPretrained, huertas-tato2024Understandingwriting] and assess authorship via similarity in the embedding space. However, such approaches remain dependent on labeled data that is often confounded with topical similarity.
The latest trend explores the use of large language models (LLMs) for human-like stylistic analysis based on prompts [hung2023WhoWrote, huang2024CanLarge]. Although promising because of their unsupervised nature, these methods require very large models in attribution settings because of context limitations, and empirical evaluation remains scarce. Our results indicate that, in practice, such approaches yield very limited performance in authorship verification.
@@ 2.4 Causal Language Modeling and few-shot learning
Causal Language Modeling (CLM) refers to conditioning on prior tokens to predict the next token in a unidirectional, autoregressive manner, ie, modeling $P(x_{t}\mid x_{<t})$ [bengio2003neuralprobabilistic]. It is the pre-training objective used by GPT-style decoder-only models [radford2019language]. A landmark discovery in large language models (LLMs) is their few‑shot in‑context learning ability [wei2022Emergentabilities]. GPT 3 (175B parameters) was shown to perform new tasks with just a few textual examples provided in the prompt, without any gradient updates or fine-tuning, demonstrating competitive performance with fully fine-tuned systems [brown2020LanguageModels]. This emergent capability highlights that massive autoregressive pre-training endows models with flexible, task‑agnostic generalization using only prompt examples.
@@ 4.1 Authorship attribution
We evaluated our models in established authorship attribution benchmarks in the fiction domain, including PAN12 [juola_2012_3713273], PAN18 [kestemont_2018_3737849], and PAN19 [kestemont_2019_3530313], and in the corporate email domain with PAN11 [argamon_2011_3713246]. Our main baseline is STAR [huertas-tato2024Understandingwriting], the strongest contrastive learning method reported in prior work. Additionally, we do not compare against LLM prompting approaches such as LIP [huang2024CanLarge], as these methods require large context windows in AA, resulting in markedly lower performance at moderate model sizes.
@@ 4.2 Authorship verification
For authorship verification (AV), we evaluate on the 2020 and 2021 editions of the PAN AV shared task [bischoff2020ImportanceSuppressing, bevendorff2020PAN20Authorship], which focus on fanfiction. In addition, we construct AV datasets from the PAN style-change detection tasks of 2022–2024 [zangerle2022PAN22Authorship, zangerle2023PAN23MultiAuthor, zangerle2024PAN24MultiAuthor]. The 2022 edition contains StackExchange posts, while the 2023 and 2024 editions comprise Reddit posts and explicitly control for topic, thereby preventing detectors from exploiting semantic information.
As in authorship attribution (AA), we compare against the STAR baseline [huertas-tato2024Understandingwriting]. Given that AV requires less context in LLM prompts than AA, we additionally include LIP [huang2024CanLarge] and PromptAV [hung2023WhoWrote] with a 24B Mistral LLM as zero-shot baselines. For STAR and OSST-based methods, we select a decision boundary with 2000 examples from the training set.
@@
The results on authorship verification datasets are summarized in figs.˜4 and 5. Several key patterns emerge: (i) Performance on challenging tasks:nOSST consistently surpasses STAR on settings that control for topical correlations (the third tasks in PAN23ST and PAN24ST), even in its zero-shot (ZS) configuration with an 8B model. In contrast, STAR maintains an advantage in datasets with strong topic biases. (ii) Comparison to LLM prompting:At comparable model scales (up to 24B parameters), nOSST methods substantially outperform LLM-based prompting approaches. (iii) Impact of the RS variant:The RS variant of nOSST generally improves performance across multiple datasets, though a few exceptions remain. (iv) Scaling behavior:Performance exhibits a clear scaling trend: larger nOSST base models tend to achieve higher accuracy, with only minor deviations from monotonic improvement.
@@ 5 Conclusions
We introduced OSST-based methods, a family of unsupervised approaches to authorship attribution and verification that leverage the extensive causal language modeling pre-training of modern LLMs. This paradigm enables scalable authorship analysis without relying on labeled data prone to topical biases.
Empirically, OSST-based methods outperform LLM prompting alternatives of comparable scale and surpass contrastively trained baselines when topical correlations are controlled. Our findings also suggest promising scaling behaviors: performance improves with model size and, for the verification random sampling variant (nOSST-RS), with the number of anchor examples, up to a saturation point. These trends indicate a practical trade-off between computational resources and stylistic accuracy.
The present work is limited by model size and by the requirement of in-distribution anchors in the RS variant, which may limit real-world applicability. Future research should explore scaling laws with larger base models and alternative anchor selection mechanisms. One such alternative could be the use of contrastively trained models for retrieval. We believe these directions can advance unsupervised stylistic modeling and bring LLM-based authorship analysis closer to robust, real-world deployment.
@@ Appendix A Model and prompt to obtain a neutrally styled version of a text
We use a google/gemma-3-12b-it [team2025Gemma3] model to obtain neutral versions of a given text. This decision was made for its multilingual capabilities and the inclusion of multilingual texts in some of the datasets used.
@@ Appendix D Base models used in OSST
  • google/gemma-3-1b-pt [team2025Gemma3]
  • google/gemma-3-4b-pt [team2025Gemma3]
  • meta-llama/Llama-3.2-1B [grattafiori2024Llama3]
  • meta-llama/Llama-3.2-8B [grattafiori2024Llama3]
@@ Appendix E Hardware and software
Appendix F Do we need neutral examples?
A natural question is whether conditioning on a neutral rewrite is actually required. To assess this, we compare our approach against a baseline (NN) that directly continues the input text without access to a neutral variant. As shown in fig.˜10, omitting the neutral version has little impact (and can even be advantageous) on datasets where topical cues are strongly aligned. However, in settings with minimal topical correlation, performance drops sharply when the task is not cast as neutralization. This is unsurprising: generating a continuation of one text conditioned on another is substantially more difficult when their topics diverge, especially in the initial tokens, which makes distinguishing related from unrelated pairs significantly easier.
Figure 10: Comparison between using (nOSST) and not using (nOSST-NN) a neutral version to compute log-probs